What is the probability that 11 out of 11 people would answer a yes/no question with yes? And how did you figure that out?
Answer by stanbon(75887)

What is the probability that 11 out of 11 people would answer a yes/no question with yes? And how did you figure that out?
-----
Each person has 1/2 a chance of saying "yes".
-----
Answer: (1/2)^11 = 1/2^11 = 0.00048828...
-----------------------------------------
Cheers,
Stan H.
=======================